One number is eight less than five times another. If the sum of the two numbers is 28, find the numbers.
Answer by Maths68(1474)

sum of the two numbers is 28
x+y=28
y=28-x.........(1)
One number is eight less than five times another
x=5y-8
Put the value of y in above equation
x=5(28-x)-8
x=140-5x-8
x+5x=132
6x=132
6x/6=132/6
x=22
Put the value of x in (1)
y=28-22
y=6
Numbers are 22, 6